My question is: The difference of 21 and the product of 5 and a number y is less than 7.
Please help ?

Answer by rfer(16322)

You can solve the problem like a regular equation then change the + to less than.
21-5y=7
-5y=7-21
5y=14
y=2.8
so y must be less than 2.8